The weights of 1,000 children, in average, is 62kg with standard deviation of 19kg. Suppose the weights are normally distributed, how many children weigh greater than 56kg?


1
Expert's answer
2022-02-07T15:41:03-0500

Solution:

Given : mean= 62

sd = 19


formula Z= "\frac{x-mean}{sd}"


For x= 56, Z= "\frac{56-62}{19}=-0.32"


Using Z table , Pr(X>56) = Pr(Z">" −0.32)=Pr(Z"\le" 0.32)=0.62552

Hence number of children = 1000"\times" 0.62552= 625.52 approximately 626 children